Sodium hydroxide (NaOH) reacts with hydrochloric acid (HCl) to form sodium chloride (NaCl) and water (H2O) according to the equation below.
Given the atomic masses: Na = 23, O = 16, and H = 1
NaOH(aq) + HCl(aq)  NaCl(aq) + H2O(l)
Which substance will be in excess if 10g of sodium hydroxide reacts with 10g of hydrochloric acid?
1
Expert's answer
2018-09-13T05:25:02-0400
m(NaOH) = 10 g
M(NaOH) = 23 + 16 +1 = 40 g/mol
n(NaOH) = m/M = 10g/40g/mol = 0.25 mol
m(HCl) = 10 g
M(HCl) = 1 + 35.5 = 36.5 g/mol
n(HCl) = 10g/36.5g/mol = 0.27 mol
According to the equation 1 mol NaOH reacts with 1 mol HCl.
HCl will be in excess.